The lumen method is to be applied to design an overhead lighting system for a room, with basic illuminance equation: total
Answer #1

2. Given

N= 500 200 2700*0.64*0.8

Floor area

Awp = 60m2

Target illuminance Ewp = 400 lux

Light loss factor = 0.8

System efficacy = 85 lm/W

power density =7.5 W/m2

To find

a) Power consumption rate at power density 7.5 W/m2

Tot =NF = System efficacy * Power density * Area

=85*7.5*60

= 38250 lumins

Power consumption == Lumin s Efficacy

38250 88

= 450 W

b) Range of coefficient of utilization

CL, 400*60 38250= C *0.8

Cu= 0.784 to 1

C) Annual cost of electricity

= 356*10*0.45*0.15

= $262.8/year